# Constraining scalar-tensor theories from higher harmonics with GW230529

### v1

@@ Abstract
The Advanced LIGO and Virgo collaborations recently detected a gravitational wave event, GW230529_181500, during the fourth observing run, which is most plausibly attributed to the merger of a neutron star and a black hole. This observation provides an opportunity to test a class of gravitational theories that deviate from general relativity. In such theories, additional terms contribute to the gravitational wave signal only in cases of asymmetric binaries. This study focuses on two scalar-tensor models within this class of theories: the Screened Modified Gravity and Brans-Dicke theory. These models have potential applications in areas such as dark matter, dark energy, cosmic inflation, and primordial nucleosynthesis. With the GW230529_181500 and Bayesian Markov-chain Monte Carlo analyses, we derive a 90% credible lower bound as $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.7\times 10^{-2}$ and $\omega_{\rm BD}>25.12$ by using dominant-mode correction. Asymmetric binary systems usually have a significant mass ratio, in such cases, higher harmonic modes cannot be neglected. Our work considers higher harmonic corrections from scalar-tensor theories and provides a tighter constraint of $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.5\times 10^{-2}$ and $\omega_{\rm BD}>32.68$, with a 13.3% and 30.1% improvement respectively. Combining GW230529_181500, GW200115 and GW190814 and including higher modes, the constraint is improved to $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<7.84\times 10^{-3}$ and $\omega_{\rm BD}>123.75$. This is currently the strongest constraint from GWs, contingent upon GW190814 being an NSBH event.
@@ 1 Introduction
In comparison to other methodologies such as laboratory and Solar System experiments, or binary pulsar and cosmological observations, GW events offer a powerful means of testing general relativity (GR) under strong-field or highly dynamical conditions [4, 5, 6, 7, 8, 9]. Such scenarios are particularly useful for distinguishing GR from modified gravity theories. As the number of detected GW events grows, they become increasingly valuable tools for placing constraints on modified gravity frameworks, such as Einstein-dilaton-Gauss-Bonnet gravity [10, 11, 12, 13, 14], extra dimension [10], time-varying gravitational constant G(t) [15] and scalar-tensor theories [10, 16, 17].
@@
Scalar-tensor theories [8, 18, 19], one of the most natural extensions of GR, introduce additional scalar degrees of freedom that couple nonminimally to the metric tensor in the gravitational sector. Such scalar fields can arise through compactification mechanisms from higher-dimensional theories like string theory [20], Kaluza-Klein-like models [21], or braneworld scenarios [22, 23]. Scalar-tensor frameworks offer a comprehensive platform to explore the phenomenology of fundamental theories and have been applied to study phenomena such as the Universe’s accelerated expansion [24, 25, 26], inflation [27, 28, 29], large-scale structure formation [30], and primordial nucleosynthesis [31, 32, 33, 34]. Here, we focus on two different models of scalar-tensor theories, i.e., the screened modified gravity (SMG) and the Brans-Dicke (BD) theory. Screened modified gravity theories, scalar degrees of freedom are introduced that can serve as dynamical dark energy components while evading stringent local gravity constraints through screening mechanisms [27]. These mechanisms, such as the Chameleon [35, 36], Vainshtein [37, 38], and Symmetron [39] mechanisms, enable the suppression of deviations from GR in high-density or small-scale environments, thereby ensuring consistency with precise laboratory and Solar System tests. On cosmological scales, however, the scalar field can remain unscreened and dynamically drive the late-time acceleration of the Universe, offering a unified framework for modified gravity and dark energy phenomenology. Extensive studies have explored the theoretical consistency, astrophysical consequences, and cosmological signatures of Screened modified gravity theories [40, 27, 41]. For Brans-Dicke theory [42], originally developed by Jordan, Fierz, Brans, and Dicke, which incorporates Mach’s principle by promoting the gravitational constant to a dynamical scalar field. The interaction strength is governed by a dimensionless coupling parameter $\omega_{\rm BD}$, with general relativity recovered in the limit $\omega_{\rm BD}\to\infty$. Extensive observational efforts have placed bounds on $\omega_{\rm BD}$ using various systems. For instance, binary pulsar timing has constrained $\omega_{\rm BD}>25000$ from the orbital decay of PSR J1738+0333 [43], while the most stringent limit to date, $\omega_{\rm BD}>40000$, arises from the Shapiro time delay measured by the Cassini mission in the Solar System [44, 7].
@@
The observational sources currently available for constraining scalar-tensor theory are largely limited to asymmetric binary systems, including NSBH, white dwarf–neutron star, and white dwarf–black hole binaries. GW230529 provides a valuable new data point for this purpose. Our work incorporates higher harmonic modes into the waveform, as such effects cannot be neglected in asymmetric binary systems with relatively large mass ratios. Specifically, we employ the open-source software PyCBC [53] and Bayesian inference methods [54, 55, 56] within the parameterized post-Einsteinian (ppE) framework [57, 58, 59, 60] to analyze GW data. Using the IMRPhenomXHM waveform model, our analysis of GW230529 data yields a constraint on the scalar-tensor coupling parameter of $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.7\times 10^{-2}$ and $\omega_{\rm BD}>25$. By including higher-mode effects we obtain constraints of $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.5\times 10^{-2}$ and $\omega_{\rm BD}>32.68$. The former represents a 13.3% and 30.1% improvement over the dominant-mode case. Meanwhile, using the observations of GW230529, GW200115, and GW190814, which is assumed to be an NSBH event, the constraints improve further, such that $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<7.84\times 10^{-3}$ and $\omega_{\rm BD}>123.75$. This is currently the strongest constraint utilizing GWs.
@@ 2.1 Parameter post-Einsteinian framework
In this paper, the GR waveform will be produced using IMRPhenomXHM, and the higher harmonics modes $(\ell,|m|)=\{(2,2),(2,1),(3,3),(3,2),(4,4)\}$ are considered. Both $\{a_{\ell m},\alpha_{\ell m},b_{\ell m},\beta_{\ell m}\}$ can be different for different values of $\ell$ and $m\,$.
We focus on the phase correction because of the negligibility of the amplitude correction. Hence the ppE waveform reduces to
@@
$\beta_{m}=\left(\frac{2}{m}\right)^{b/3-1}\beta_{22},$ (2.5)
@@ 2.2 Waveform in scalar-tensor theory
Here, $\tilde{g}^{\mu\nu}$ and $\tilde{g}$ denote the Einstein-frame metric and its determinant, respectively. $\tilde{R}\equiv\tilde{g}^{\mu\nu}\tilde{R}^{\mu\nu}$ is the Ricci scalar. $\varphi$ denotes the scalar field, and $V(\varphi)$ is its potential, which gives the field an effective mass. For Brans–Dicke theory, the potential term $V(\varphi)$ is not considered. $\Psi$ denotes the matter fields and $\phi(\varphi)$ is the conformal coupling function. In scalar-tensor theories, compact binary systems exhibit a misalignment between the centers of gravitational binding energy and inertial mass. This asymmetry leads to a time-varying dipolar structure, which radiates additional energy through scalar emission [65]. In our analysis, we adopt a waveform model incorporating the leading-order correction to general relativity, characterized by a dipole contribution to the gravitational-wave phase [65, 60, 66, 67],
@@
For screened modified gravity theory, we only consider the leading order modification and replace $\alpha_{\rm A}-\alpha_{\rm B}$ as $\varphi_{\rm{}_{VEV}}/({M_{\rm Pl}\Phi_{A}})$. Where $M_{\rm Pl}=\sqrt{1/8\pi G}$ is the reduced Planck mass, $\varphi_{\rm VEV}$ is the vacuum expectation value of the scalar field, and $\Phi_{A}=Gm/R$ is the surface gravitational potential of the object $A$. In conjunction with Eq.(2.1), the dominant-mode correction of these ppE parameters in screened modified gravity theory can be written as [43, 60, 67]:
@@
For Brans-Dicke theory, where one can replace $(\alpha_{\rm A}-\alpha_{\rm B})^{2}$ as $2(s_{\rm A}-s_{\rm B})^{2}/(2+\omega_{\rm BD})$ [43]. Here, $s_{\rm A}$ denotes the sensitivity of the $A$th body, which approximately equals its compactness (about 0.5 for black holes and $\sim 0.2$ for neutron stars). Combining the insights from Eq.(2.1), the dominant-mode correction of these ppE parameters in BD theory can be written as [43, 60, 67]:
@@
The leading order modification in the GW waveform introduced by scalar-tensor theory starts at the PN order, corresponding to $b^{\rm SMG}=b^{\rm BD}=-7$. And by combining Eqs.(2.4) and Eq.(2.5), we can derive $\beta_{\ell m}^{\rm SMG}$ and $\beta_{\ell m}^{\rm BD}$.
@@ 3 Bayesian inference
where $p(\bm{\vartheta}|\mathcal{H})$ represents the prior distribution over parameters under the hypothesis $\mathcal{H}$, and $p(d|\bm{\vartheta},\mathcal{H})$ is the likelihood function, which quantifies the probability of observing the data $d$ assuming the parameters $\bm{\vartheta}$. The likelihood is not derived directly from the waveform model itself, but rather constructed by statistically comparing the observed data with theoretical waveform predictions—such as those generated under the parameterized post-Einsteinian (ppE) formalism—while accounting for the noise properties of the detectors. Typically, the noise is modeled as stationary and Gaussian. The denominator $p(d|\mathcal{H})$ is known as the Bayesian evidence, which serves as a normalization factor and encapsulates the overall plausibility of the hypothesis $\mathcal{H}$ given the data.
@@
[…]
where $S_{n}(f)$ denotes the one-sided power spectral density (PSD) of the detector noise, and $(\tilde{a}(f),\tilde{b}(f),\tilde{a}^{*}(f),\tilde{b}^{*}(f))$ are the Fourier transforms of the time-domain signals $a(t)$ and $b(t)$, respectively. $f_{\rm low}$ denotes the low-frequency cutoff of the gravitational wave data, $f_{\rm max}=f_{\rm ISCO}=(6^{3/2}\pi M)^{-1}$ represents the upper cutoff frequency of inspiral stage.
@@ 4 Results
Figure 1: Probability density of Bayesian posterior distributions on $\varphi_{-2}$. The gray line corresponds to the constraint at a 90% credible level by using the dominant mode (DM) correction, while the red line stands for the constraint at the 90% confidence level by using the correction including the higher mode (HM).
The analysis is carried out using the open-source PyCBC software package [53], employing MCMC sampling with the emcee_pt sampler [71], initialized with 200 walkers. To mitigate contamination from low-frequency noise, we adopt a lower frequency cutoff of $f_{\rm low}=20$ Hz. A data segment of 64 seconds centered around GW230529 is used for parameter estimation. As for prior choices, we assume a uniform distribution for $1/(\omega_{\rm BD}+2)$ in the range $[0,0.5]$, and for $\varphi_{-2}$ in the interval $[-0.0027,0]$. Spin components $\chi_{1z}$ and $\chi_{2z}$ are aligned or anti-aligned with the orbital angular momentum, and are independently drawn from a uniform distribution in the range $[-0.99,0.99]$.
Figure 2: Probability density of Bayesian posterior distributions on $\omega_{\rm{}_{BD}}$. The gray line corresponds to the constraint on BD at 90% credible level by using the dominant mode (DM) correction, while the red line represents the constraint on BD with 90% probability by using the correction including higher mode (HM). We use a logarithmic scale due to the posterior distribution of $\omega_{\rm BD}$ spanning several orders of magnitude.
@@
Figure 3: Probability density of Bayesian posterior distributions on $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}$. The gray line corresponds to the constraint at a 90% credible level by using the dominant-mode (DM) correction, while the red line stands for the constraint at the 90% confidence level by using the correction including the higher mode (HM).
@@
We present the resulting constraints on $\varphi_{-2}$ in scalar–tensor gravity in Fig. 1. For GW230529, we obtain a 90% credible lower bound of $\varphi_{-2}>-1.98\times 10^{-4}$ using the dominant mode (DM) correction only. When higher-order modes HM are included, the constraint improves to $\varphi_{-2}>-1.54\times 10^{-4}$, corresponding to a 22.22% enhancement. Corresponding constraints on the Brans–Dicke coupling parameter $\omega_{\rm BD}$ are shown in Fig. 2. For GW230529, the dominant-mode analysis yields a 90% lower bound of $\omega_{\rm BD}>25$. Including higher-order modes improves the bound to $\omega_{\rm BD}>32.68$, a 30.1% enhancement. For sreened modified gravity theory, we adopt neutron star parameters $m=1.4M_{\odot}$ and $R=10$km. Using the same Bayesian pipeline and the GW230529 signal, we constrain the scalar field vacuum expectation value $\varphi_{\rm VEV}$. We find that $\varphi_{\rm VEV}/M_{\rm Pl}<1.7\times 10^{-2}$ using the dominant mode, and $\varphi_{\rm VEV}/M_{\rm Pl}<1.5\times 10^{-2}$ when higher-order modes are included.
@@
We also consider a combined analysis of the three NSBH events, GW230529, GW200115 and GW190814. Due to convergence issues in the dominant-mode-only case [51], we report only the result using higher-order modes for the combined dataset: $\varphi_{-2}>-4.26\times 10^{-5}$, demonstrating substantial improvement. In the combined analysis of the three NSBH events incorporating higher modes, the constraint strengthens significantly to $\omega_{\rm BD}>123.75$. Similarly, Considering the combination, we obtain a tighter constraint as well $\varphi_{\rm VEV}/M_{\rm Pl}<7.84\times 10^{-3}$. This represents the most stringent bound to date using gravitational-wave observations, assuming GW190814 is indeed an NSBH event. All the above resulting bounds are summarized in Table 2.
@@
$\varphi_{-2}$ $-1.98\times 10^{-4}$ $-1.54\times 10^{-4}$ $-4.26\times 10^{-5}$

### v2

@@ Abstract
The Advanced LIGO and Virgo collaborations recently detected a gravitational wave event, GW230529_181500, during the fourth observing run, which is most plausibly attributed to the merger of a neutron star and a black hole. This observation provides an opportunity to test a class of gravitational theories that deviate from general relativity. In such theories, additional terms contribute to the gravitational wave signal only in cases of asymmetric binaries. This study focuses on two scalar-tensor models within this class of theories: the Screened Modified Gravity and Brans-Dicke theory. These models have potential applications in areas such as dark matter, dark energy, cosmic inflation, and primordial nucleosynthesis. With the GW230529_181500 and Bayesian Markov-chain Monte Carlo analyses, we derive a 90% credible lower bound as $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.7\times 10^{-2}$ and $\omega_{\rm BD}>25.12$ by using dominant mode correction. Asymmetric binary systems usually have a significant mass ratio, in such cases, higher harmonic modes cannot be neglected. Our work considers higher harmonic corrections from scalar-tensor theories and provides a tighter constraint of $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.5\times 10^{-2}$ and $\omega_{\rm BD}>32.68$, with a 13.3% and 30.1% improvement respectively. Combining GW230529_181500, GW200115 and GW190814 and including higher modes, the constraint is improved to $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<7.84\times 10^{-3}$ and $\omega_{\rm BD}>123.75$. This is currently the strongest constraint from GWs, contingent upon GW190814 being an NSBH event.
@@ 1 Introduction
In comparison to other methodologies such as laboratory and Solar System experiments, or binary pulsar and cosmological observations, GW events offer a powerful means of testing general relativity (GR) under strong-field or highly dynamical conditions [4, 5, 6, 7, 8, 9]. Such scenarios are particularly useful for distinguishing GR from modified gravity theories. As the number of detected GW events grows, they become increasingly valuable tools for placing constraints on modified gravity frameworks, such as Einstein–scalar–Gauss–Bonnet (including the dilaton case, often referred to as Einstein–dilaton–Gauss–Bonnet) gravity [10, 11, 12, 13, 14, 15, 16], extra dimension [11], time-varying gravitational constant G(t) [17] and scalar-tensor theories [11, 18, 19, 20, 21, 22, 23, 24, 25, 26].
@@
Scalar-tensor theories [8], one of the most natural extensions of GR, introduce additional scalar degrees of freedom that couple nonminimally to the metric tensor in the gravitational sector. Such scalar fields can arise through compactification mechanisms from higher-dimensional theories like string theory [27], Kaluza-Klein-like models [28], or braneworld scenarios [29, 30]. Scalar-tensor frameworks offer a comprehensive platform to explore the phenomenology of fundamental theories and have been applied to study phenomena such as the Universe’s accelerated expansion [31, 32, 33], inflation [34, 35, 36], large-scale structure formation [37], and primordial nucleosynthesis [38, 39, 40, 41]. Here, we focus on two different models of scalar-tensor theories, i.e., the screened modified gravity (SMG) and the Brans-Dicke (BD) theory. Screened modified gravity theories, scalar degrees of freedom are introduced that can serve as dynamical dark energy components while evading stringent local gravity constraints through screening mechanisms [34]. These mechanisms, such as the Chameleon [42, 43], Vainshtein [44, 45], and Symmetron [46] mechanisms, enable the suppression of deviations from GR in high-density or small-scale environments, thereby ensuring consistency with precise laboratory and Solar System tests. On cosmological scales, however, the scalar field can remain unscreened and dynamically drive the late-time acceleration of the Universe, offering a unified framework for modified gravity and dark energy phenomenology. Extensive studies have explored the theoretical consistency, astrophysical consequences, and cosmological signatures of Screened modified gravity theories [47, 34, 48, 49]. For Brans-Dicke theory [50], originally developed by Jordan, Fierz, Brans, and Dicke, which incorporates Mach’s principle by promoting the gravitational constant to a dynamical scalar field. The interaction strength is governed by a dimensionless coupling parameter $\omega_{\rm BD}$, with general relativity recovered in the limit $\omega_{\rm BD}\to\infty$. Extensive observational efforts have placed bounds on $\omega_{\rm BD}$ using various systems. For instance, binary pulsar timing has constrained $\omega_{\rm BD}>25000$ from the orbital decay of PSR J1738+0333 [51], while the most stringent limit to date, $\omega_{\rm BD}>40000$, arises from the Shapiro time delay measured by the Cassini mission in the Solar System [52, 7].
@@
The observational sources currently available for constraining scalar-tensor theory are largely limited to asymmetric binary systems, including NSBH, white dwarf–neutron star, and white dwarf–black hole binaries. GW230529 provides a valuable new data point for this purpose. Our work incorporates higher harmonic modes into the waveform, as such effects cannot be neglected in asymmetric binary systems with relatively large mass ratios. Specifically, we employ the open-source software PyCBC [61] and Bayesian inference methods [62, 63, 64] within the parameterized post-Einsteinian (ppE) framework [65, 66, 67, 68] to analyze GW data. Using the IMRPhenomXHM waveform model, our analysis of GW230529 data yields a constraint on the scalar-tensor coupling parameter of $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.7\times 10^{-2}$ and $\omega_{\rm BD}>25$. By including higher-mode effects we obtain constraints of $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<1.5\times 10^{-2}$ and $\omega_{\rm BD}>32.68$. The former represents a 13.3% and 30.1% improvement over the dominant mode case. Meanwhile, using the observations of GW230529, GW200115, and GW190814, which is assumed to be an NSBH event, the constraints improve further, such that $\frac{\varphi_{\rm{}_{VEV}}}{M_{\rm Pl}}<7.84\times 10^{-3}$ and $\omega_{\rm BD}>123.75$. This is currently the strongest constraint utilizing GWs.
@@ 2.1 Parameter post-Einsteinian framework
In this paper, the GR waveform will be produced using IMRPhenomXHM, and the higher harmonics modes $(\ell,|m|)=\{(2,2),(2,1),(3,3),(3,2),(4,4)\}$ are considered. Both $\{a_{\ell m},\alpha_{\ell m},b_{\ell m},\beta_{\ell m}\}$ can be different for different values of $\ell$ and mm\,.
We focus on phase corrections, since Tahura et al. [12] validated the common approach of considering only phase corrections. They showed that including amplitude corrections together with phase corrections changes the inferred bounds by at most $\sim 4\%$ across several modified gravity theories. Thus, following the standard practice in the literature, we adopt the phase-only ppE waveform, which reduces to
@@
$\beta_{m}=\left(\frac{2}{m}\right)^{b/3-1}\beta_{22}.$ (2.5)
It is worth noting that this differs from the waveform treatment adopted by Sänger et al., who performed a theory agnostic analysis by introducing generic deformations, and by construction their prescription corresponds within the ppE framework to $\beta_{\ell m}=(m/2)\,\beta_{22}$ [10]. In addition, this also differs from the waveform adopted by Gao et al., which treats $\beta_{\ell m}$ uniformly across modes [16].
@@ 2.2 Waveform in scalar-tensor theory
Here, $\tilde{g}^{\mu\nu}$ and $\tilde{g}$ denote the Einstein-frame metric and its determinant, respectively. $\tilde{R}\equiv\tilde{g}^{\mu\nu}\tilde{R}^{\mu\nu}$ is the Ricci scalar. $\varphi$ denotes the scalar field, and $V(\varphi)$ is its potential, which gives the field an effective mass. For Brans–Dicke theory, the potential term $V(\varphi)$ is not considered. $\Psi$ denotes the matter fields and $\phi(\varphi)$ is the conformal coupling function. In scalar-tensor theories, compact binary systems exhibit a misalignment between the centers of gravitational binding energy and inertial mass. This asymmetry leads to a time-varying dipolar structure, which radiates additional energy through scalar emission [73]. In our analysis, we adopt a waveform model incorporating the leading order correction to general relativity, characterized by a dipole contribution to the gravitational wave phase for the dominant mode [73, 68, 74, 75],
@@
For screened modified gravity theory, we only consider the leading order modification and replace $\alpha_{\rm A}-\alpha_{\rm B}$ as $\varphi_{\rm{}_{VEV}}/({M_{\rm Pl}\Phi_{A}})$. Where $M_{\rm Pl}=\sqrt{1/8\pi G}$ is the reduced Planck mass, $\varphi_{\rm VEV}$ is the vacuum expectation value of the scalar field, and $\Phi_{A}=Gm/R$ is the surface gravitational potential of the object $A$. In conjunction with Eq.(2.1), the dominant mode correction of these ppE parameters in screened modified gravity theory can be written as [51, 68, 75]:
@@
For Brans-Dicke theory, where one can replace $(\alpha_{\rm A}-\alpha_{\rm B})^{2}$ as $2(s_{\rm A}-s_{\rm B})^{2}/(2+\omega_{\rm BD})$ [51]. Here, $s_{\rm A}$ denotes the sensitivity of the $A$th body, which approximately equals its compactness (about 0.5 for black holes and $\sim 0.2$ for neutron stars). Combining the insights from Eq.(2.1), the dominant mode correction of these ppE parameters in BD theory can be written as [51, 68, 75]:
@@
The leading order modification in the GW waveform introduced by scalar-tensor theory starts at the 1-1PN order, corresponding to $b^{\rm SMG}=b^{\rm BD}=-7$. And by combining Eqs.(2.4) and Eq.(2.5), we can derive $\beta_{\ell m}^{\rm SMG}$ and $\beta_{\ell m}^{\rm BD}$.
@@ 3 Bayesian inference
where $p(\bm{\vartheta}|\mathcal{H})$ represents the prior distribution over parameters under the hypothesis $\mathcal{H}$, and $p(d|\bm{\vartheta},\mathcal{H})$ is the likelihood function, which quantifies the probability of observing the data $d$ assuming the parameters $\bm{\vartheta}$. The likelihood is not derived directly from the waveform model itself, but rather constructed by statistically comparing the observed data with theoretical waveform predictions, such as those generated under the parameterized post-Einsteinian (ppE) formalism while accounting for the noise properties of the detectors. Typically, the noise is modeled as stationary and Gaussian. The denominator $p(d|\mathcal{H})$ is known as the Bayesian evidence, which serves as a normalization factor and encapsulates the overall plausibility of the hypothesis $\mathcal{H}$ given the data.
@@
$\left<a(t)\middle|b(t)\right>=2\int_{f_{\rm low}}^{f_{\rm high}}\frac{\tilde{a}^{*}(f)\tilde{b}(f)+\tilde{a}(f)\tilde{b}^{*}(f)}{S_{n}(f)}df,$ (3.3)
where $S_{n}(f)$ denotes the one-sided power spectral density (PSD) of the detector noise, and $(\tilde{a}(f),\tilde{b}(f),\tilde{a}^{*}(f),\tilde{b}^{*}(f))$ are the Fourier transforms of the time-domain signals $a(t)$ and $b(t)$, respectively. $f_{\rm low}$ denotes the low-frequency cutoff of the gravitational wave data, $Mf_{\mathrm{max}}=0.018$ represents the intermediate transition frequency of the inspiral stage, as specified by the phenomenological model [79, 80, 81].
@@ 4 Results
Figure 1: Probability density of Bayesian posterior distributions on $\varphi_{-2}$. The blue dashed line corresponds to the 90% credible constraint using the dominant mode correction, while the red solid line denotes the 90% credible constraint using the correction including higher modes. The gray dash-dotted line represents the prior on $\varphi_{-2}$.
The analysis is carried out using the open-source PyCBC software package [61], employing MCMC sampling with the emcee_pt sampler [82], initialized with 200 walkers. To mitigate contamination from low-frequency noise, we adopt a lower frequency cutoff of $f_{\rm low}=20$ Hz. A data segment of 64 seconds centered around GW230529 is used for parameter estimation. As for prior choices, we assume a uniform distribution for $1/(\omega_{\rm BD}+2)$ in the range $[0,0.5]$, and for $\varphi_{-2}$ in the interval [0.0027,0][-0.0027,0] (Some theory-agnostic analyses adopt priors spanning both positive and negative values; for a comparison with negative-only priors and the resulting impact on the posterior, see Appendix A). Spin components $\chi_{1z}$ and $\chi_{2z}$ are aligned or anti-aligned with the orbital angular momentum, and are independently drawn from a uniform distribution in the range [0.99,0.99][-0.99,0.99].
Figure 2: Probability density of Bayesian posterior distributions on $\omega_{\rm{}_{BD}}$. The blue dashed line corresponds to the constraint on BD at 90% credible level by using the dominant mode correction, while the red solid line represents the constraint on BD with 90% probability by using the correction including higher mode. The gray dash-dotted line represents the prior on $\omega_{\rm{}_{BD}}$. We use a logarithmic scale due to the posterior distribution of $\omega_{\rm BD}$ spanning several orders of magnitude.
@@
Figure 3: Probability density of Bayesian posterior distributions on ${\varphi_{\rm{}_{VEV}}}/{M_{\rm Pl}}$. The blue dashed line corresponds to the constraint at a 90% credible level by using the dominant mode correction, while the red solid line stands for the constraint at the 90% confidence level by using the correction including the higher mode. The gray dash-dotted line represents the prior on ${\varphi_{\rm{}_{VEV}}}/{M_{\rm Pl}}$.
@@
We present the resulting constraints on $\varphi_{-2}$ in scalar–tensor gravity in Fig. 1. For GW230529, we obtain a 90% credible lower bound of $\varphi_{-2}>-1.98\times 10^{-4}$ using the dominant mode (DM) correction only. It is consistent, in order of magnitude, with Sänger et al.’s ESGB constraint based on the dominant mode only [10]. When higher-order modes HM are included, the constraint improves to $\varphi_{-2}>-1.54\times 10^{-4}$, corresponding to a 22.22% enhancement. Corresponding constraints on the Brans–Dicke coupling parameter $\omega_{\rm BD}$ are shown in Fig. 2. For GW230529, the dominant mode analysis yields a 90% lower bound of $\omega_{\rm BD}>25$. Including higher-order modes improves the bound to $\omega_{\rm BD}>32.68$, a 30.1% enhancement. For sreened modified gravity theory, we adopt neutron star parameters $m=1.4M_{\odot}$ and R=10R=10km. Using the same Bayesian pipeline and the GW230529 signal, we constrain the scalar field vacuum expectation value $\varphi_{\rm VEV}$. We find that $\varphi_{\rm VEV}/M_{\rm Pl}<1.7\times 10^{-2}$ using the dominant mode, and $\varphi_{\rm VEV}/M_{\rm Pl}<1.5\times 10^{-2}$ when higher-order modes are included.
[…]
For Vainshtein-type screening [45, 85], the suppression of scalar charges is controlled by the Vainshtein radius, $r_{V}\sim\left(\frac{GM}{\Lambda^{3}}\right)^{1/3}$, as in the cubic Galileon case, where $\Lambda$ denotes the strong-coupling scale. For typical parameter values considered in Galileon and massive gravity models, compact objects are well within their Vainshtein radii, so dipole radiation is highly suppressed. Our constraint therefore limits the $\varphi_{\rm VEV}$, but it does not yield a unique bound on $r_{V}$ without committing to a specific model. For K-essence screening (kinetic screening/k-mouflage) [86, 87], the effective scalar charge depends on the functional form of the kinetic term $K(X)$ and its derivatives. While our constraint excludes large bare couplings, a quantitative mapping to a “screening radius” or scalar sound speed again requires specifying the detailed form of $K(X)$. In this sense, our result is generic in ruling out sizable effective scalar charges at the orbital scales probed by GW230529, across different screening mechanisms. However, translating this bound into a constraint on a screening radius or other model-specific parameter is only possible once a concrete Lagrangian is chosen.
We also consider a combined analysis of the three NSBH events, GW230529, GW200115 and GW190814. For the stacking procedure [88, 89, 14], we follow the methodology described in Perkins et al. [14]. Specifically, we implemented both of the approaches outlined there. In the first method, for each event we construct a histogram of the MCMC posterior samples and then obtain a smooth analytic representation of the posterior by applying kernel density estimation (KDE). We then multiply the resulting KDEs pointwise across events to form the joint posterior. In the second, more direct method, we align the binning across events, construct histograms of the posterior samples, and then multiply these histograms bin by bin, which is equivalent to multiplying the marginalized likelihoods of independent events. In both cases, mild smoothing is applied when evaluating confidence intervals in order to mitigate numerical fluctuations at the distribution tails. Throughout this work we quote the more conservative of the two results.
@@
Due to convergence issues in the dominant mode only case for GW190814 [59], We cannot report the combined result for this using the dominant mode. Therefore, we report only the result using higher-order modes for the combined dataset: $\varphi_{-2}>-4.26\times 10^{-5}$, demonstrating substantial improvement. We follow the same principle, and quote the more conservative constraint of the two methods. In the combined analysis of the three NSBH events incorporating higher modes, the constraint strengthens significantly to $\omega_{\rm BD}>123.75$. Similarly, Considering the combination, we obtain a tighter constraint as well $\varphi_{\rm VEV}/M_{\rm Pl}<7.84\times 10^{-3}$. All the above resulting bounds are summarized in Table 2.
@@
$\varphi_{-2}$ $1.98\times 10^{-4}$ $1.54\times 10^{-4}$ $4.26\times 10^{-5}$